OpenStudy (anonymous):

36 inches is 12% of what length?

OpenStudy (anonymous):

would it be 300 in

OpenStudy (senkai):

Yes, it would be 300 inches.

OpenStudy (senkai):

Pretty sure anyway, haha

OpenStudy (senkai):

Cause 300 x .12 = 36

OpenStudy (anonymous):

36 = 12/100x (36 is equal to twelve one hundredths of some number that we'll call x. 12/100 is 12 percent) Solve by multiplying both sides by the inverse of the fraction which is 100/12. x = 36 * 100/12 x = 3600/12 x = 300